### v1

@@ 1 INTRODUCTION
Kilonovae are transients powered by the radioactive decay of heavy nuclei synthesized by the rapid neutron-capture process ($r$-process) in ejecta from binary neutron star mergers (BNSM) or black hole-neutron star (BH-NS) mergers. For the first BNSM event detected in gravitational waves by Advanced LIGO and Virgo, GW170817 (Abbott et al., 2017a), the electromagnetic counterparts have also been observed from radio to gamma-rays (Abbott et al., 2017b; Alexander et al., 2017; Chornock et al., 2017; Hallinan et al., 2017; McCully et al., 2017; Savchenko et al., 2017; Ruan et al., 2018; Balasubramanian et al., 2021; Hajela et al., 2022). For the kilonova counterpart, several follow-up optical and infrared searches recorded detailed photometric and spectroscopic data for $\sim 10$ days after the merger, though late-time detections in certain bands continued for weeks to months (Abbott et al., 2017b; Kasen et al., 2017; Pian et al., 2017). Most theoretical kilonova models suggest that GW170817 BNSM ejecta possibly consists of two major components – a lanthanide-poor component dominating the earlier phase of the blue emission, and a lanthanide-rich component accounting for the later transition to the infrared (e.g. McCully et al. 2017; Kasen et al. 2017; Metzger 2020; Margutti & Chornock 2021).
@@
Current numerical simulations of BNSMs have found several mass ejection mechanisms in the dynamical phase of the merger, as well as during the secular evolution of the post-merger remnants (e.g. Rosswog et al. 1999; Dessart et al. 2008; Bauswein et al. 2013; Hotokezaka et al. 2013; Fernández & Metzger 2013; Shibata et al. 2017; Nedora et al. 2019; Shibata & Hotokezaka 2019.) Predicting properties relevant to kilonova modeling, such as the ejecta mass, velocity and the nucleosynthesis yields still depends on various physical inputs that are currently not fully known. These include the equation of state that determines the post-merger lifetime of the hypermassive neutron star (HMNS) (e.g. Bauswein et al. 2013; Hotokezaka et al. 2013; Lucca & Sagunski 2020; Kawaguchi et al. 2023), the treatment of viscosity arising from the turbulent magnetic fields (e.g. Fernández et al. 2019; Kiuchi et al. 2018; Siegel & Metzger 2018; Ruiz et al. 2019; Kawaguchi et al. 2022), the modeling of neutrino-matter interaction and neutrino flavor oscillations (e.g. Radice et al. 2016; Joy & Rosch 2022; Li & Siegel 2021; Just et al. 2022; Fernández et al. 2022; Foucart 2023; Espino et al. 2024), or the potential ejecta-ejecta or jet-ejecta interaction (e.g. Geng et al. 2019; Hamidani et al. 2020; Nativi et al. 2021; Gottlieb et al. 2022), among others. While current predictions from BNSM simulations are broadly consistent with observations (Metzger, 2020; Margutti & Chornock, 2021), much more efforts remain for quantitatively accurate kilonova modeling.
@@
Among the various ejecta components, the post-merger disk outflow is expected to dominate given the inferred parameters for GW170817 (e.g., Shibata et al. 2017). One of the key remaining uncertainties in predicting ejecta from disk outflows is the impact of $r$-process heating. Most long-term post-merger (magneto-)hydrodynamical simulations include energy release from nuclear recombination of alpha particles (4He) and/or from heavier nuclei whose abundances are determined by the nuclear statistical equilibrium (NSE), which works together with viscous heating driving the expansion of the disk and the associated mass outflow (Fernández & Metzger, 2013; Just et al., 2015; Kawaguchi et al., 2021). Using a large number ($\sim 4,000$) of heavy nuclei with temperature-dependent partition functions in NSE can result in an additional 1.7 MeV per nucleon released relative to using only one representative heavy nucleus, which can have a non-negligible impact on the outflow velocity (Fernández et al., 2024). However, this treatment does not take into account the additional energy release due to the $r$-process when non-equilibrium nuclear reactions drive further changes in nuclear compositions when the temperature drops below $\sim 6$ GK. While including the energy release due to nuclear transmutations beyond NSE by directly coupling a full nuclear reaction network with multidimensional (magneto-)hydrodynamical simulations is beyond reach, several works have attempted to address the potential impact of additional nuclear heating on properties of different ejecta components from BNSMs and BH-NS mergers as well as the fallback material (Metzger et al., 2010a; Rosswog et al., 2014; Fernández et al., 2015; Just et al., 2015; Wu et al., 2016; Desai et al., 2019; Ishizaki et al., 2021; Klion et al., 2022; Foucart et al., 2021; Darbha et al., 2021; Kawaguchi et al., 2021; Sneppen et al., 2023; Kawaguchi et al., 2024; Magistrelli et al., 2024; Just et al., 2025). These works have adopted very different ways of approximating the nuclear heating as an additional source term. They found different levels of impact due to nuclear heating beyond NSE on ejecta mass and velocity, and their spatial distributions that can possibly affect the detailed modeling of kilonova light curves (Rosswog et al., 2014; Just et al., 2015; Wu et al., 2016; Klion et al., 2022; Foucart et al., 2021; Darbha et al., 2021; Kawaguchi et al., 2021; Sneppen et al., 2023; Kawaguchi et al., 2024; Magistrelli et al., 2024), as well as the amount and timescale of fallback matter that may be linked to the observation of $\gamma$-ray bursts (Metzger et al., 2010a; Fernández et al., 2015; Desai et al., 2019; Ishizaki et al., 2021; Foucart et al., 2021), highlighting the need of incorporating $r$-process heating in simulations.
@@
Particularly for the disk outflows, earlier investigations in Just et al. (2015) and Wu et al. (2016) implemented an additional nuclear heating rate that depends only on the local temperature and the sign of the radial velocity of the fluid in axisymmetric, viscous hydrodynamic simulations for BH-disk winds. While the temperature-dependent heating rates adopted in Just et al. (2015) and Wu et al. (2016) are based on post-processed nuclear reaction network calculations of tracer particles from simulations without including $r$-process heating, they neglected the strong dependence of heating rates on nuclear composition. The impact of additional nuclear heating on ejecta mass was found to be potentially significant in Wu et al. (2016) but negligible in Just et al. (2015), possibly related to the different amount of included additional heating at $T\gtrsim 2-3$ GK. Klion et al. (2022) implemented a simple time-dependent heating parametrization, also assumed to be composition-independent, in long-term hydrodynamical simulations mapped from a three-dimensional general relativistic magneto-hydrodynamical BH-disk simulation conducted for 10 s. They found that $r$-process heating increases the ejecta velocity by about a factor of 2 when reaching the phase of homologous expansion at $\sim 60$ s, which results in a brighter and bluer kilonova. Similar to Klion et al. (2022), Kawaguchi et al. (2021) and Kawaguchi et al. (2024) included $r$-process heating in long-term hydrodynamical simulations for disk outflows mapped from numerical relativity simulations of BNSMs after a few seconds, where the central remnants are massive NSs. On top of that, they partially took into account the composition and ejection time dependence of $r$-process heating in long-term simulations by including the nuclear heating rates obtained in post-processed nucleosynthesis calculations based on tracer particles of the numerical relativity simulations, and found relatively minor impact on ejecta mass and velocity.
@@
In this study, we also aim to investigate the potential impact of $r$-process heating on the evolution of the post-merger BH-disk system and its outflow properties by including its composition and spatial/temporal dependence. We assume that the composition dependence of the nuclear heating rate mainly depends on the electron fraction ($Y_{e}$) at $T\simeq 6$ GK. We utilize a large number of “memory” tracer particles, an externally-tabulated $Y_{e}$ dependent heating rate, and the technique of Cloud-in-Cell (CIC) method (Birdsall & Fuss, 1969) to distribute the heating rate information carried by the memory particles to the simulation domain. In Section 2, we describe the numerical methods adopted for our hydrodynamic simulations, including the developed heating method. In Section 3, we present the overall evolution observed in our simulations. Sections 4, 5 and 6 discuss the main results, focusing on the details of $r$-process heating feedback. Finally, we summarize our findings in Section 8.
2 METHOD AND MODELS
2.1 Hydrodynamic Simulation
@@
We use the FLASH code, version 3.2 (Fryxell et al., 2000; Dubey et al., 2009), to perform viscous hydrodynamic simulations with the setup described in Fernández & Metzger (2013); Metzger & Fernández (2014); Fernández et al. (2015), and Fernández et al. (2022). The simulations employ two-dimensional axisymmetric spherical coordinates. The computational domain extends radially from $\sim$ 8.8 km to $8.8\times 10^{5}$ km, logarithmically divided into 640 cells, resulting in a cell size of $\Delta r\sim 1.49$ km for the innermost cell, with outflow boundaries at both ends. In the polar direction, the angle $\theta$ spans from 0 to $\pi$, distributed evenly in terms of $\cos{\theta}$ across 112 cells, employing reflecting boundaries. The finest resolution at the equator corresponds to $\Delta\theta\sim 0.018$ radian. The code employs domain decomposition to parallelize execution using the Message Passing Interface (e.g., Clarke et al. 1994), with each processor working with a domain block of equal number of cells.
@@
At the center of the simulation box, we place a spinning BH with a mass of 2.65 $M_{\odot}$ and spin parameter of $0.8$ as the remnant compact object after the merger binary, which acts on its surroundings through a pseudo-Newtonian gravitational potential (Artemova et al., 1996). The simulations are initialized from an equilibrium torus characterized by constant specific angular momentum, a constant entropy of 8 $k_{\rm B}$/baryon, and an electron fraction ($Y_{e}$) of 0.1. The disk has a total mass of 0.1 $M_{\odot}$, with the maximum density occurring at a radius of 50 km. We implement an $\alpha$-disk model, having an azimuthal viscous stress with a viscosity coefficient $\alpha=0.03$ or 0.06, which quantifies the strength of angular momentum transport (Shakura & Sunyaev, 1973).
@@
In our simulations, we use the Helmholtz equation of state (Timmes & Swesty, 2000), with protons, neutrons, and alpha particles in nuclear statistical equilibrium. The nuclear binding energy of alpha particles is included in the internal energy. A 3-species leakage scheme with emission and annular lightbulb-type absorption is employed for the neutrino transport, with the updates described in Fernández et al. (2022). Passive tracer particles are used to trace the fluid properties throughout the whole simulation. In each simulation, we deploy $10^{4}$ tracers with equal effective mass, initially distributed according to the density profile (cf. Wu et al. 2016).
2.2 Implementation of $r$-process heating feedback
@@
To achieve the goal of implementing the composition-dependent nuclear heating rate during $r$-process, we adopt the rate of nuclear energy release profiles from nucleosynthesis calculations performed over a parametrized expansion trajectory with an assumed entropy per nucleon of $30$ $k_{\rm B}$ and an expansion time scale $\tau=10$ ms (Lippuner & Roberts, 2015; Wu et al., 2019) for 50 different initial $Y_{e}$ (at 8 GK) values uniformly spaced from 0.01 to 0.50. Figure 1 shows the nuclear energy release rates as functions of temperature for all 50 different initial $Y_{e}$, highlighting extended (shortened) nuclear energy release for lower (higher) $Y_{e}$ values. For simplicity, we assume that 50% of the released nuclear energy is carried away by neutrinos, with the remaining half contributing to the fluid’s internal energy as extra heating source (Metzger et al., 2010b; Barnes et al., 2016).
@@
At each time step, we map the fluid temperature and $Y_{e}$ onto each tracer. When a tracer passes through the temperature of 6 GK for the first time, its $Y_{e}$ is recorded and designated as the initial $Y_{e}$ in our heating profile; we denote this value as $Y_{e,T6}$. Heating is applied to the outflow material in our fiducial models only when the local fluid temperature drops below a threshold temperature ($T_{\rm heat}$) of 4 GK during the simulation, approximating the onset of the $r$-process. To assess the sensitivity of this threshold, we also consider a higher threshold temperature of 6 GK (see Section 5). We apply the same heating rate with a negative sign for inflowing material ($v_{r}<0$), which serves as cooling. This treatment prevents convecting tracers from contributing redundant heating energy when they flow outward multiple times. The specific heating rate is calculated through interpolation based on temperature and $Y_{e,T6}$ for each tracer. The interpolation procedure involves three steps: (1) identifying the closest two tabulated $Y_{e,T6}$ values that bracket the tracer’s $Y_{e,T6}$; (2) performing temperature interpolation in log scale at each of these $Y_{e,T6}$ values using the tracer’s temperature; and (3) interpolating between the two resulting heating rates in linear scale to account for differences in $Y_{e,T6}$. For tracers with $Y_{e,T6}>0.5$, we adopt the heating rate corresponding to $Y_{e,T6}=0.5$ to avoid extrapolation beyond the boundary of the table. The computed specific heating rate is then subsequently mapped back onto the grid-based fluid elements. Besides taking different threshold temperature, we also present a resolution test for the adopted number of $Y_{e}$ bins used for heating rate interpolation by introducing a low-resolution model, al03T4Ye5, which uses 5 bins instead of 50. See Appendix B.1 for detail discussion.
Throughout this study, we employ the Cloud-in-Cell (CIC) method (Birdsall & Fuss, 1969) to map information between grid cells and tracers. In the mesh-to-particle mapping, field quantities such as temperature and $Y_{e}$ are assigned to each tracer using a first-order distance-weighted interpolation based on the values of the cell containing the tracer and its adjacent cells. Conversely, in the particle-to-mesh mapping, tracer properties (the specific heating rate) are distributed to the surrounding grid cells using the same distance-weighted scheme, where each tracer’s contribution to a given cell is proportional to its distance to the cell center. This method ensures smooth transitions across cell boundaries and allows a limited number of tracers to exchange information with a larger set of grid cells.
2.3 Models

### v2

@@ I INTRODUCTION
Kilonovae are transients powered by the radioactive decay of heavy nuclei synthesized by the rapid neutron-capture process ($r$-process) in ejecta from binary neutron star mergers (BNSM) or black hole-neutron star (BH-NS) mergers. For the first BNSM event detected in gravitational waves by Advanced LIGO and Virgo, GW170817 (Abbott et al., 2017a), the electromagnetic counterparts have also been observed from radio to gamma-rays (Abbott et al., 2017b; Alexander et al., 2017; Chornock et al., 2017; Hallinan et al., 2017; McCully et al., 2017; Savchenko et al., 2017; Ruan et al., 2018; Balasubramanian et al., 2021; Hajela et al., 2022). For the kilonova counterpart, several follow-up optical and infrared searches recorded detailed photometric and spectroscopic data for $\sim 10$ days after the merger, though late-time detections in certain bands continued for weeks to months (Abbott et al., 2017b; Kasen et al., 2017; Pian et al., 2017). Most theoretical kilonova models suggest that GW170817 BNSM ejecta possibly consists of two major components – a lanthanide-poor component dominating the earlier phase of the blue emission, and a lanthanide-rich component accounting for the later transition to the infrared (e.g. McCully et al. 2017; Kasen et al. 2017; Metzger 2020; Margutti and Chornock 2021).
@@
Current numerical simulations of BNSMs have found several mass ejection mechanisms in the dynamical phase of the merger, as well as during the secular evolution of the post-merger remnants (e.g. Rosswog et al. 1999; Dessart et al. 2008; Bauswein et al. 2013; Hotokezaka et al. 2013; Fernández and Metzger 2013; Shibata et al. 2017b; Nedora et al. 2019; Shibata and Hotokezaka 2019.) Predicting properties relevant to kilonova modeling, such as the ejecta mass, velocity and the nucleosynthesis yields still depends on various physical inputs that are currently not fully known. These include the equation of state that determines the post-merger lifetime of the hypermassive neutron star (HMNS) (e.g. Bauswein et al. 2013; Hotokezaka et al. 2013; Lucca and Sagunski 2020; Kawaguchi et al. 2023), the treatment of viscosity arising from the turbulent magnetic fields (e.g. Fernández et al. 2019; Kiuchi et al. 2018; Siegel and Metzger 2018; Ruiz et al. 2019; Kawaguchi et al. 2022), the modeling of neutrino-matter interaction and neutrino flavor oscillations (e.g. Radice et al. 2016; Joy and Rosch 2022; Li and Siegel 2021; Just et al. 2022; Fernández et al. 2022; Foucart 2023; Espino et al. 2024), or the potential ejecta-ejecta or jet-ejecta interaction (e.g. Geng et al. 2019; Hamidani et al. 2020; Nativi et al. 2021; Gottlieb et al. 2022), among others. While current predictions from BNSM simulations are broadly consistent with observations (Metzger, 2020; Margutti and Chornock, 2021), much more efforts remain for quantitatively accurate kilonova modeling.
@@
Among the various ejecta components, the post-merger disk outflow is expected to dominate given the inferred parameters for GW170817 (e.g., Shibata et al. 2017a). One of the key remaining uncertainties in predicting ejecta from disk outflows is the impact of $r$-process heating. Most long-term post-merger (magneto-)hydrodynamical simulations include energy release from nuclear recombination of alpha particles (4He) and/or from heavier nuclei whose abundances are determined by the nuclear statistical equilibrium (NSE), which works together with viscous heating driving the expansion of the disk and the associated mass outflow (Fernández and Metzger, 2013; Just et al., 2015; Kawaguchi et al., 2021). Using a large number ($\sim 4,000$) of heavy nuclei with temperature-dependent partition functions in NSE can result in an additional 1.7 MeV per nucleon released relative to using only one representative heavy nucleus, which can have a non-negligible impact on the outflow velocity (Fernández et al., 2024). However, this treatment does not take into account the additional energy release due to the $r$-process when non-equilibrium nuclear reactions drive further changes in nuclear compositions when the temperature drops below $\sim 6$ GK. While including the energy release due to nuclear transmutations beyond NSE by directly coupling a full nuclear reaction network with multidimensional (magneto-)hydrodynamical simulations is beyond reach, several works have attempted to address the potential impact of additional nuclear heating on properties of different ejecta components from BNSMs and BH-NS mergers as well as the fallback material (Metzger et al., 2010a; Rosswog et al., 2014; Fernández et al., 2015b; Just et al., 2015; Wu et al., 2016; Desai et al., 2019; Ishizaki et al., 2021; Klion et al., 2022; Foucart et al., 2021; Darbha et al., 2021; Kawaguchi et al., 2021; Sneppen et al., 2023; Kawaguchi et al., 2024; Magistrelli et al., 2024; Just et al., 2025). These works have adopted very different ways of approximating the nuclear heating as an additional source term. They found different levels of impact due to nuclear heating beyond NSE on ejecta mass and velocity, and their spatial distributions that can possibly affect the detailed modeling of kilonova light curves (Rosswog et al., 2014; Just et al., 2015; Wu et al., 2016; Klion et al., 2022; Foucart et al., 2021; Darbha et al., 2021; Kawaguchi et al., 2021; Sneppen et al., 2023; Kawaguchi et al., 2024; Magistrelli et al., 2024), as well as the amount and timescale of fallback matter that may be linked to the observation of $\gamma$-ray bursts (Metzger et al., 2010a; Fernández et al., 2015b; Desai et al., 2019; Ishizaki et al., 2021; Foucart et al., 2021), highlighting the need of incorporating $r$-process heating in simulations.
@@
Particularly for the disk outflows, earlier investigations in Just et al. (2015) and Wu et al. (2016) implemented an additional nuclear heating rate that depends only on the local temperature and the sign of the radial velocity of the fluid in axisymmetric, viscous hydrodynamic simulations for BH-disk winds. While the temperature-dependent heating rates adopted in Just et al. (2015) and Wu et al. (2016) are based on post-processed nuclear reaction network calculations of tracer particles from simulations without including $r$-process heating, they neglected the strong dependence of heating rates on nuclear composition. The impact of additional nuclear heating on ejecta mass was found to be potentially significant in Wu et al. (2016) but negligible in Just et al. (2015), possibly related to the different amount of included additional heating at $T\gtrsim 2-3$ GK. Klion et al. (2022) implemented a simple time-dependent heating parametrization, also assumed to be composition-independent, in long-term hydrodynamical simulations mapped from a three-dimensional general relativistic magneto-hydrodynamical BH-disk simulation conducted for 10 s. They found that $r$-process heating increases the ejecta velocity by about a factor of 2 when reaching the phase of homologous expansion at $\sim 60$ s, which results in a brighter and bluer kilonova. Similar to Klion et al. (2022), Kawaguchi et al. (2021) included r-process heating in long-term hydrodynamical simulations for disk outflows mapped from numerical relativity simulations of BNSMs with massive neutron star remnants. In a related study, Kawaguchi et al. (2024) performed analogous simulations but for BH-NS mergers, where the central remnant is a black hole. Both studies partially took into account the composition and ejection time dependence of r-process heating in long-term simulations by including the nuclear heating rates obtained from post-processed nucleosynthesis calculations based on tracer particles from numerical relativity simulations. They found relatively minor impacts on ejecta mass and velocity.
@@
In this study, we also aim to investigate the potential impact of $r$-process heating on the evolution of the post-merger BH-disk system and its outflow properties by including its composition and spatial/temporal dependence. We assume that the composition dependence of the nuclear heating rate mainly depends on the electron fraction ($Y_{e}$) at $T\simeq 6$ GK. We utilize a large number of “memory” tracer particles, an externally-tabulated $Y_{e}$ dependent heating rate, and the technique of Cloud-in-Cell (CIC) method (Birdsall and Fuss, 1969) to distribute the heating rate information carried by the memory particles to the simulation domain. In Section II, we describe the numerical methods adopted for our hydrodynamic simulations, including the developed heating method. In Section III, we present the overall evolution observed in our simulations. Sections IV, V and VI discuss the main results, focusing on the details of $r$-process heating feedback. Finally, we summarize our findings in Section VIII.
II METHOD AND MODELS
II.1 Hydrodynamic Simulation
@@
We use the FLASH code, version 3.2 (Fryxell et al., 2000; Dubey et al., 2009), to perform viscous hydrodynamic simulations with the setup described in Fernández and Metzger (2013); Metzger and Fernández (2014); Fernández et al. (2015a), and Fernández et al. (2022). The simulations employ two-dimensional axisymmetric spherical coordinates. The computational domain extends radially from $\sim$ 8.8 km to $8.8\times 10^{5}$ km, logarithmically divided into 640 cells, resulting in a cell size of $\Delta r\sim 1.49$ km for the innermost cell, with outflow boundaries at both ends. In the polar direction, the angle $\theta$ spans from 0 to $\pi$, distributed evenly in terms of $\cos{\theta}$ across 112 cells, employing reflecting boundaries. The finest resolution at the equator corresponds to $\Delta\theta\sim 0.018$ radian. The code employs domain decomposition to parallelize execution using the Message Passing Interface (e.g., Clarke et al. 1994), with each processor working with a domain block of equal number of cells.
@@
The initial condition is set to match the inferred parameters of the event GW170817 with a post-merger BH remnant and an accretion disk (Abbott et al., 2017a; Fernández et al., 2022). At the center of the simulation box, we place a spinning BH with a mass of 2.65 $M_{\odot}$ and spin parameter of $0.8$ as the remnant compact object after the merger binary, which acts on its surroundings through a pseudo-Newtonian gravitational potential (Artemova et al., 1996). The simulations are initialized from an equilibrium torus characterized by constant specific angular momentum, a constant entropy of 8 $k_{\rm B}$/baryon, and an electron fraction ($Y_{e}$) of 0.1. The disk has a total mass of 0.1 $M_{\odot}$, with the maximum density occurring at a radius of 50 km. We implement an $\alpha$-disk model, having an azimuthal viscous stress with a viscosity coefficient $\alpha=0.03$ or 0.06, which quantifies the strength of angular momentum transport (Shakura and Sunyaev, 1973).
@@
In our simulations, we use the Helmholtz equation of state (Timmes and Swesty, 2000), with protons, neutrons, and alpha particles in nuclear statistical equilibrium. The nuclear binding energy of alpha particles is included in the internal energy. A 3-species leakage scheme with emission and annular lightbulb-type absorption is employed for the neutrino transport, with the updates described in Fernández et al. (2022). Passive tracer particles are used to trace the fluid properties throughout the whole simulation. In each simulation, we deploy $10^{4}$ tracers with equal effective mass, initially distributed according to the density profile (cf. Wu et al. 2016).
II.2 Implementation of $r$-process heating feedback
@@
To achieve the goal of implementing the composition-dependent nuclear heating rate during $r$-process, we adopt the rate of nuclear energy release profiles from nucleosynthesis calculations performed over a parametrized expansion trajectory with an assumed entropy per nucleon of $30$ $k_{\rm B}$ and an expansion time scale $\tau=10$ ms (Lippuner and Roberts, 2015; Wu et al., 2019) for 50 different initial $Y_{e}$ (at 8 GK) values uniformly spaced from 0.01 to 0.50. The specific parameters, $s=30$ $k_{\rm B}$ per nucleon and $\tau=10$ ms, were chosen because the $\rho(T)$ curve resulting from this profile is representative of the thermal trajectories observed in the simulated outflow tracers. Figure 1 shows the nuclear energy release rates as functions of temperature for all 50 different initial $Y_{e}$, highlighting extended (shortened) nuclear energy release for lower (higher) $Y_{e}$ values. For simplicity, we assume that 50% of the released nuclear energy is carried away by neutrinos, with the remaining half contributing to the fluid’s internal energy as extra heating source (Metzger et al., 2010b; Barnes et al., 2016).
@@
At each time step, we map the fluid temperature and $Y_{e}$ onto each tracer. When a tracer passes through the temperature of 6 GK for the first time, its $Y_{e}$ is recorded and designated as the initial $Y_{e}$ in our heating profile; we denote this value as $Y_{e,T6}$. Heating is applied to the outflow material in our fiducial models only when the local fluid temperature drops below a threshold temperature ($T_{\rm heat}$) of 4 GK during the simulation, approximating the onset of the $r$-process. To assess the sensitivity of this threshold, we also consider a higher threshold temperature of 6 GK (see Section V). We apply the same heating rate with a negative sign for inflowing material ($v_{r}<0$), which serves as cooling. This treatment prevents convecting tracers from contributing redundant heating energy when they flow outward multiple times. The specific heating rate is calculated through interpolation based on temperature and $Y_{e,T6}$ for each tracer. The interpolation procedure involves three steps: (1) identifying the closest two tabulated $Y_{e,T6}$ values that bracket the tracer’s $Y_{e,T6}$; (2) performing temperature interpolation in log scale at each of these $Y_{e,T6}$ values using the tracer’s temperature; and (3) interpolating between the two resulting heating rates in linear scale to account for differences in $Y_{e,T6}$. For tracers with $Y_{e,T6}>0.5$, we adopt the heating rate corresponding to $Y_{e,T6}=0.5$ to avoid extrapolation beyond the boundary of the table. The computed specific heating rate is then subsequently mapped back onto the grid-based fluid elements. Besides taking different threshold temperature, we also present a resolution test for the adopted number of $Y_{e}$ bins used for heating rate interpolation by introducing a low-resolution model, al03T4Ye5, which uses 5 bins instead of 50. See Appendix B.1 for detail discussion.
Throughout this study, we employ the Cloud-in-Cell (CIC) method (Birdsall and Fuss, 1969) to map information between grid cells and tracers. In the mesh-to-particle mapping, field quantities such as temperature and $Y_{e}$ are assigned to each tracer using a first-order distance-weighted interpolation based on the values of the cell containing the tracer and its adjacent cells. Conversely, in the particle-to-mesh mapping, tracer properties (the specific heating rate) are distributed to the surrounding grid cells using the same distance-weighted scheme, where each tracer’s contribution to a given cell is proportional to its distance to the cell center. This method ensures smooth transitions across cell boundaries and allows a limited number of tracers to exchange information with a larger set of grid cells.
II.3 Models